THE ROTATIONAL SPECTRA OF THE CARBON CHAIN ANIONS C_{2n}H^{−} (n = 1, 2, 4)

\underline{SANDRA BRÜNKEN}, CARL A. GOTTLIEB, HARSHAL GUPTA^{a}, MICHAEL C. McCARTHY, AND PATRICK THADDEUS, Harvard Smithsonian Center for Astrophysics, and Division of Engineering and Applied Sciences, Harvard University, Cambridge, MA 02138, USA.

Our recent discovery that the carbon chain anion C_{6}H^{−} is the carrier of a series of hitherto unidentified rotational transitions towards the carbon-rich circumstellar shell IRC+10216^{b}, and the detection of this anion in a second standard molecular source, TMC-1^{c}, prompted us to investigate the related anions CCH^{−}, C_{4}H^{−}, and C_{8}H^{−} in the laboratory. Here, we present highly accurate gas-phase measurements of the rotational spectrum of these closed shell anions^{d}^{e}. Lines of CCH^{−} and C_{4}H^{−} were observed in the millimeter-wave band with a free-space millimeter-wave spectrometer in a dc discharge of acetylene and argon. The rotational spectrum of C_{8}H^{−}, and additional transitions of C_{4}H^{−}, were recorded in the centimeter-wave band with our molecular beam Fourier transform microwave (FTM) spectrometer. The anions were produced in a discharge of diacetylene heavily diluted in an inert buffer gas. The derived spectroscopic constants allow for the precise determination of rotational line frequencies for astronomical searches throughout the radio band. Based on this work, the C_{4}H^{−} anion has now been detected in IRC+10216^{f}. The status of a search for C_{8}H^{−} in space will be reported.

aalso at: Institute for Theoretical Chemistry, Department of Chemistry and Biochemistry, The University of Texas, Austin, TX 78712, USA

bK. Kawaguchi, Y. Kasai, S.-I. Ishikawa, and N. Kaifu, PASJ 47 (1995), 853

cM. C. McCarthy, C. A. Gottlieb, H. Gupta, and P. Thaddeus, ApJL 652 (2006), L141

dS. Brünken, C. Gottlieb, H. Gupta, M. C. McCarthy, and P. Thaddeus, A&AL, in press (2007)

eH. Gupta, S. Brünken, F. Tamassia, C. A. Gottlieb, M. C. McCarthy, and P. Thaddeus, ApJL, 655 (2007), L57

fM. Guélin, J. Cernicharo, M. Agundez, M. C. McCarthy, P. Thaddeus, and K. Kawaguchi, submitted to ApJL (2007)